Ridgewood Mills is a division of Iowa Woolen Products. For the most recent year, Ridgewood had net income of $20,000,000. Included in income was interest expense of $1,450,000. The operation’s tax rate is 40 percent. Total assets of Ridgewood Mills are $230,000,000, current liabilities are $52,400,000, and $36,000,000 of the current liabilities are noninterest bearing.

Calculate NOPAT, invested capital, and ROI for Ridgewood Mills. (Round ROI to 2 decimal places, e.g. 15.25%.)

NOPAT $

Invested capital $

ROI %
Answer #1

NOPAT = Net income+(Interest expense after tax) = 20000000+(1450000*60%) = 20870000

Invested capital = 230000000-52400000 = 177600000

ROI = 20870000/177600000 = 11.75%